# Forecasting-Based Biomedical Time-series Data Synthesis for Open Data and Robust AI

### v1

@@ Abstract
The limited data availability due to strict privacy regulations and significant resource demands severely constrains biomedical time-series AI development, which creates a critical gap between data requirements and accessibility. Synthetic data generation presents a promising solution by producing artificial datasets that maintain the statistical properties of real biomedical time-series data without compromising patient confidentiality. We propose a framework for synthetic biomedical time-series data generation based on advanced forecasting models that accurately replicates complex electrophysiological signals such as EEG and EMG with high fidelity. These synthetic datasets preserve essential temporal and spectral properties of real data, which enables robust analysis while effectively addressing data scarcity and privacy challenges. Our evaluations across multiple subjects demonstrate that the generated synthetic data can serve as an effective substitute for real data and also significantly boost AI model performance. The approach maintains critical biomedical features while provides high scalability for various applications and integrates seamlessly into open-source repositories, substantially expanding resources for AI-driven biomedical research.
@@ 1 Introduction
Figure 1: Open-source contribution and data gap fulfillment. (a) Open-source data contribution: Synthetic datasets are uploaded to public repositories (Hugging Face, Kaggle, AWS), enabling broader access for AI development while maintaining privacy compliance. (b) Data gap fulfillment: Synthetic samples populate underrepresented regions in the feature space of real datasets, enhancing classifier training by providing a more comprehensive representation of the ideal data distribution.
@@
The development of high-performance AI models for biomedical time-series applications requires extensive and diverse datasets [1, 2]. Biomedical data availability, however, remains severely constrained due to stringent privacy regulations, substantial acquisition costs, and the rarity of certain biomedical conditions [3]. These constraints create a considerable disparity between the data necessary for robust AI development and the data currently accessible to researchers [4]. Synthetic data generation offers a powerful strategy for addressing these limitations through the creation of artificial datasets that maintain the statistical properties of real biomedical data without compromising patient confidentiality [5]. This approach significantly advances biomedical AI research by directly mitigating critical data availability challenges that currently impede progress in AI-driven applications [6].
@@
Biomedical signals exhibit intricate temporal patterns that must be accurately preserved in synthetic data to ensure biomedical validity and practical utility [7]. The time-series forecaster methods excel at understanding the temporal characteristics of sequential data better than any other models [8]. These techniques identify and replicate underlying data patterns to generate realistic continuations of time signals, preserving essential temporal properties vital for diagnostic and monitoring applications [9]. Notably, this methodology demonstrates exceptional adaptability across diverse biomedical signal types, including electromyography (EMG), electroencephalograms (EEGs), and continuous glucose monitoring data, each with unique temporal signatures that forecasting methods can effectively capture and reproduce. The resulting synthetic data derived from time-series forecasting maintains high biomedical relevance and accuracy required for meaningful applications.
@@
The integration of synthetic biomedical data into open-source repositories substantially accelerates research and innovation in biomedical-focused AI [4]. Many current open-source biomedical datasets lack sufficient size and diversity, severely restricting the development of robust and generalizable AI models. Contributing high-quality synthetic datasets to prominent platforms such as Hugging Face, Kaggle, AWS, and data.gov significantly expands available resources while democratizing data access across the biomedical AI research community. Synthetic datasets inherently protect patient privacy by generating artificial samples without personally identifiable information, while adherence to international privacy standards, including the General Data Protection Regulation (GDPR) [10], enables unrestricted global data sharing and eliminates risks of patient re-identification [11].
@@
In addition, absolute data scarcity represents a critical challenge for biomedical AI model development, particularly for rare diseases or emerging biomedical conditions where existing real-world datasets frequently lack sufficient data points to train robust and reliable AI models [12]. We address these critical shortages through strategic deployment of synthetic data generation to produce additional relevant samples that mirror the characteristics of limited real datasets [13, 14]. The AI models trained on these expanded datasets consistently demonstrate improved predictive performance and enhanced generalization capabilities [15]. This targeted synthetic data supplementation directly alleviates the issue of insufficient biomedical data availability, substantially advancing the effectiveness and applicability of AI systems in biomedical environments [16].
@@
Here, we introduce a paradigm shift approach that repurposes time-series forecasters—used for next-step prediction—as synthesizers to address the persistent challenges of biomedical AI: privacy and data scarcity. To validate our approach, we synthesized EEG sleep-stage signals using 16 state-of-the-art time-series forecasting models [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. When combined, synthetic data consistently improved performance across all subjects, with DLinear achieving a 3.71 percentage point gain. Moreover, synthetic-only training slightly outperformed original data, with SOFTS achieving 91.00% compared to 90.83%. These results confirm that forecasting-based data synthesis is a practical, scalable solution that enables privacy-preserving open data sharing and enhances AI model robustness by fulfilling underrepresented regions.
The remainder of this paper is organized as follows. Section 2 details data acquisition and the proposed forecasting-based pipeline. Section 3 reports subject-wise O/S/O+S results, window-size analyses across 16 forecasters, UMAP comparisons of original vs. synthetic data, and a benchmark against a representative GAN for time-series data. Section 4 discusses implications, limitations, and guidelines for choosing the synthetic-to-original ratio. Finally, Section 5 concludes the paper.
@@ 2.1 Animal Data Acquisition
All in vivo data was acquired from a previous study (IACUC approval: KA2021-066) [33]. Briefly, EEG/EMG signals were recorded at a sampling rate of 1 kHz using a biopotential acquisition device (RHD2000, Intan Technologies, CA, USA), which was then amplified and digitally filtered (low-pass filter at 0.1 Hz, high-pass filter at 7.5 kHz, and notch filter at 60 Hz). Filtered EEG/EMG signals were segmented into 5s epochs for further analysis using custom-written software (MATLAB, MathWorks, Natick, MA, USA). Implementation of robust fairness protocols requires systematic approaches to bias detection and mitigation throughout the generation pipeline. The development of specialized fairness constraints helps guide synthetic data generation while maintaining biomedical utility. Furthermore, regular evaluation of population representation metrics will enable continuous monitoring and adjustment of generation parameters to support equitable coverage.
Figure 2: Framework for synthetic biomedical time-series data generation. Synthetic data generation using time-series forecasters: A forecasting model trained on real biomedical signals generates extended patterns that produce synthetic datasets.
@@ 2.2 Synthetic Data Generation
Synthetic signals were generated from real EEG recordings using a class-conditional time-series forecasting approach. Let $D=\{(y^{(i)},c^{(i)})\}_{i=1}^{N}$ denote the original dataset, where $y^{(i)}\in\mathbb{R}^{T}$ represents an EEG time-series of length $T$, and $c^{(i)}\in\mathcal{C}$ is the corresponding sleep stage label from the class set $\mathcal{C}=\{\text{``WAKE''},\text{``NREM''},\text{``REM''}\}$. For each class $c\in\mathcal{C}$, a separate time-series forecasting model $f_{c}(\cdot;\theta_{c})$ was trained using supervised input-output pairs generated with a sliding window of size $L$, forecasting a future segment of fixed length $H=500$. Specifically, training pairs were constructed as:
$(x_{t},y_{t})=\big(y^{(i)}_{t:t+L-1},\;y^{(i)}_{t+L:t+L+H-1}\big)$, (1)
where $x_{t}\in\mathbb{R}^{L}$ serves as the context window and $y_{t}\in\mathbb{R}^{H}$ is the target to be predicted. For all experiments, we varied $L\in\{10,25,50,100,250\}$. The objective was to minimize the empirical Huber loss [34] over the class-specific training set:
@@
where $D_{c}=\{(y^{(i)},c^{(i)})\in D\mid c^{(i)}=c\}$ and $\mathcal{L}_{\text{Huber}}(\cdot)$ denotes the standard Huber loss. All models were trained using a batch size of 32 and a maximum of 1,000 optimization steps, with a constant sampling frequency of 100 Hz (10 ms time intervals) for uniform time indices.
After training, each model $f_{c}$ was used to generate synthetic EEG signals via recursive sliding-window prediction. Starting from an initial context window $x^{(i)}$, the model iteratively predicted chunks of length $H$, where previous synthetic outputs were used as context for subsequent predictions until the target sequence length of 500 time points was reached. The synthetic samples:
$\hat{y}^{(i)}=f^{\text{Recursive}}_{c^{(i)}}\big(x^{(i)};\theta_{c^{(i)}}\big),$ (3)
were paired with their corresponding class labels to form the final synthetic dataset $\hat{D}=\{(\hat{y}^{(i)},c^{(i)})\}_{i=1}^{N}$. Overall procedure is described in Algorithm 1.
@@
4   Build $D_{c}=\{(x_{t},y_{t})\}$ via sliding windows
5 $\theta_{c}^{\ast}\leftarrow\arg\min_{\theta}\ \frac{1}{|D_{c}|}\sum_{(x_{t},y_{t})\in D_{c}}\mathcal{L}_{\text{Huber}}\!\big(f_{c}(x_{t};\theta),\,y_{t}\big)$
@@ 11 $\hat{y}\leftarrow$[\,]
12 while $|\hat{y}|<500$ do
13    $\tilde{y}\leftarrow f_{c^{(i)}}(x^{(i)};\theta_{c^{(i)}}^{\ast})$
14      Append $\tilde{y}$ to $\hat{y}$ and update $x^{(i)}$ with the most recent $L$ points
@@ 2.3 Evaluation Framework
To evaluate synthetic data utility, we designed a comprehensive classification framework with three training conditions: original data only (O), synthetic data only (S), and combined original and synthetic data (O+S). Let $D_{\text{orig}}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}$ and $D_{\text{syn}}=\{(\hat{x}^{(i)},\hat{y}^{(i)})\}_{i=1}^{N}$ represent the original and synthetic datasets. In the O+S condition, the training set was constructed as:
@@
Each EEG time-series $x^{(i)}\in\mathbb{R}^{T}$ was converted into a time-frequency representation using the short-time Fourier transform (STFT):
$S^{(i)}(f,t)=\left|\sum_{\tau=0}^{T-1}x^{(i)}(\tau)w(\tau-t)e^{-j2\pi f\tau}\right|^{2},$ (5)
@@
Model training was performed on a fixed train/test split, with training and testing sets predefined for each subject. The classifier model was a ResNet-18 [35], trained for a maximum of 500 epochs using stochastic gradient descent with a learning rate of $10^{-4}$. Note that, while ResNet-18 served as the default classifier, the framework is compatible with other image classification backbones. To ensure reproducibility, random seeds were fixed for all training runs, and experiments were repeated with 5 different seeds.
@@
$\mathcal{L}(\theta)=-\sum_{i}\sum_{c=1}^{C}\mathbbm{1}[y^{(i)}=c]\log p_{\theta}^{(i)}(c),$ (6)
where $\mathbbm{1}[\cdot]$ is the indicator function, $p_{\theta}^{(i)}(c)$ represents the predicted probability for class $c$ on sample $i$ and $C$ is the number of classes. The original-only (O) and synthetic-only (S) training conditions followed identical preprocessing, cross-validation, and evaluation procedures as the O+S condition, ensuring consistent experimental methodology across all settings.
@@ 2.4 Performance Metrics
Model performance was quantitatively assessed using accuracy, precision, recall, and F1-score across all training conditions. Let $y^{(i)}\in\mathcal{C}$ and $\hat{y}^{(i)}\in\mathcal{C}$ denote the ground-truth and predicted class labels for the $i$-th sample in the test set, where $\mathcal{C}=\{\text{WAKE},\text{NREM},\text{REM}\}$. The primary evaluation metric was classification accuracy, computed as:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[y^{(i)}=\hat{y}^{(i)}],$ (7)
where $N$ represents the total number of test samples.
To evaluate class-wise performance, we computed precision, recall, and F1-score for each class $c\in\mathcal{C}$. Let $\text{TP}_{c}$, $\text{FP}_{c}$, and $\text{FN}_{c}$ denote the number of true positives, false positives, and false negatives for class $c$. Precision and recall were defined as:
$\text{Precision}_{c}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}},\quad\text{Recall}_{c}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FN}_{c}}.$ (8)
@@
$\text{F1}_{c}=2\cdot\frac{\text{Precision}_{c}\cdot\text{Recall}_{c}}{\text{Precision}_{c}+\text{Recall}_{c}}.$ (9)

### v2

@@ Abstract
The limited data availability due to strict privacy regulations and significant resource demands severely constrains biomedical time-series AI development, which creates a critical gap between data requirements and accessibility. Synthetic data generation presents a promising solution by producing artificial datasets that maintain the statistical properties of real biomedical time-series data without compromising patient confidentiality. While GANs, VAEs, and diffusion models capture global data distributions, forecasting models offer inductive biases tailored for sequential dynamics. We propose a framework for synthetic biomedical time-series data generation based on recent forecasting models that accurately replicates complex electrophysiological signals such as EEG and EMG with high fidelity. These synthetic datasets can be freely shared for open AI development and consistently improve downstream model performance. Numerical results on sleep-stage classification show up to a 3.71% performance gain with augmentation and a 91.00% synthetic-only accuracy that surpasses the real-data-only baseline.
@@ 1 Introduction
Figure 1: Open-source contribution and data gap fulfillment. (a) Open-source data contribution: Synthetic datasets are uploaded to public repositories (Hugging Face, Kaggle, AWS), enabling broader access for open AI development while maintaining privacy compliance. (b) Data gap fulfillment: Synthetic samples populate underrepresented regions in the feature space of real datasets, enhancing classifier training by providing a more comprehensive representation of the ideal data distribution.
@@
The development of high-performance AI models for biomedical time-series applications requires extensive and diverse datasets [acosta2022multimodal, zhou2023foundation]. Biomedical data availability, however, remains severely constrained due to stringent privacy regulations, substantial acquisition costs, and the rarity of certain biomedical conditions [moor2023foundation]. These constraints create a considerable disparity between the data necessary for robust AI development and the data currently accessible to researchers [mahmood2025benchmarking]. Synthetic data generation offers a powerful strategy for addressing these limitations through the creation of artificial datasets that maintain the statistical properties of real biomedical data without compromising patient confidentiality [chen2021synthetic]. This approach significantly advances biomedical AI research by directly mitigating critical data availability challenges that currently impede progress in AI-driven applications [van2024synthetic].
@@
Biomedical signals contain complex inter-signal patterns that must be retained in synthetic data to ensure biomedical relevance and practical utility [norori2021addressing]. Moreover, the variability of temporal dynamics across signal types poses additional challenges for synthetic data generation. The time-series forecaster methods excel at understanding the temporal characteristics of sequential data better than any other models [torres2021deep]. These techniques identify and replicate underlying data patterns to generate realistic continuations of time signals, preserving essential temporal properties vital for diagnostic and monitoring applications [chen2023long]. Notably, this methodology demonstrates exceptional adaptability across diverse biomedical signal types, including electromyography (EMG), electroencephalograms (EEGs), and continuous glucose monitoring data, each with unique temporal signatures that forecasting methods can effectively capture and reproduce. Thus, the resulting synthetic data derived from time-series forecasting maintains high biomedical relevance and accuracy required for meaningful applications.
@@
The integration of synthetic biomedical data into open-source repositories substantially accelerates research and innovation in biomedical-focused AI [mahmood2025benchmarking]. Many current open-source biomedical datasets lack sufficient size and diversity, severely restricting the development of robust and generalizable AI models. Contributing high-quality synthetic datasets to prominent platforms such as Hugging Face, Kaggle, AWS, and data.gov significantly expands available resources while democratizing data access across the biomedical AI research community. Moreover, the synthetic datasets inherently protect patient privacy by generating artificial samples without personally identifiable information, while adherence to international privacy standards, including the General Data Protection Regulation (GDPR) [regulation2018general], enables unrestricted global data sharing and eliminates risks of patient re-identification [yoon2020anonymization].
@@
In addition, absolute data scarcity represents a critical challenge for biomedical AI model development, particularly for rare diseases or emerging biomedical conditions where existing real-world datasets frequently lack sufficient data points to train robust and reliable AI models [bansal2022systematic]. We address these critical shortages through strategic deployment of synthetic data generation to produce additional relevant samples that mirror the characteristics of limited real datasets [chen2024towards, loecher2021using]. The AI models trained on these expanded datasets consistently demonstrate improved predictive performance and enhanced generalization capabilities [PEZOULAS20242892]. This synthetic data supplementation directly alleviates the issue of insufficient biomedical data availability, substantially advancing the effectiveness and applicability of AI systems in biomedical environments [gao2023synthetic].
@@
Despite the strengths of GANs [yoon2019time, esteban2017real], VAEs [desai2021timevae, pmlr-v206-lee23d], and diffusion models [tashiro2021csdi, timeweaver2024] in general data synthesis, the sequential nature of biomedical time signals requires models that explicitly incorporate temporal inductive biases. Here, we introduce a paradigm-shifting approach that repurposes time-series forecasters—used for next-step prediction—as synthesizers to address the persistent challenges of biomedical AI: privacy and data scarcity (Fig. 1). By leveraging forecasting models, our approach generates high-fidelity sequences that better capture real-world signal dynamics than general-purpose baselines. To validate our approach, we synthesized EEG sleep-stage signals using 16 state-of-the-art time-series forecasting models [dialtedrnn, tcn, nbeats, deepar, tft, bitcn, nbeatsx, nhits, dlinear, patchtst, timesnet, tide, deepnpts, itransformer, softs, kan]. When combined, synthetic data consistently improved performance across all subjects, with DLinear achieving a 3.71 percentage point gain. Moreover, synthetic-only training slightly outperformed original data, with SOFTS achieving 91.00% compared to 90.83%. The results indicate that forecasting-based synthesis is a practical approach that supports privacy-preserving data sharing and improves model robustness by enriching underrepresented regions.
The remainder of this paper is organized as follows. Section 2 details the animal data acquisition procedure, the forecasting-based synthesis pipeline, and the evaluation framework. Section 3 reports the experimental findings, including data preparation, performance and class-wise analyses, visualization of original versus synthetic data, quantitative quality and privacy assessments, and comparison with representative GAN-based synthesis. Section 4 analyzes the effect of the synthetic-to-original data ratio on performance, highlights key limitations, and proposes directions for future work. Finally, Section 5 provides concluding remarks.
@@ 2.1 Animal Data Acquisition
All in vivo data was acquired from a previous study (IACUC [^1] Institutional Animal Care and Use Committee approval: KA2021-066) [jo2022general]. Briefly, EEG/EMG signals were recorded at a sampling rate of 1 kHz using a biopotential acquisition device (RHD2000, Intan Technologies, CA, USA), which was then amplified and digitally filtered (low-pass filter at 0.1 Hz, high-pass filter at 7.5 kHz, and notch filter at 60 Hz). Filtered EEG/EMG signals were segmented into 5s epochs for further analysis using custom-written software (MATLAB, MathWorks, Natick, MA, USA).
Figure 2: Framework for synthetic biomedical time-series data generation. Synthetic data generation using time-series forecasters: The forecasting model trained on real biomedical signals generates similar patterns or enriches underrepresented segments.
@@ 2.2 Synthetic Data Generation
The synthetic signals were generated from real EEG recordings using a class-conditional time-series forecasting approach (Fig. 2). Formally, let $D=\{(y^{(i)},c^{(i)})\}_{i=1}^{N}$ denote the original dataset, where $y^{(i)}$ represents an EEG time-series, and $c^{(i)}\in\mathcal{C}$ is the corresponding sleep stage label from the class set $\mathcal{C}=\{\text{``WAKE''},\text{``NREM''},\text{``REM''}\}$.
For each class $c\in\mathcal{C}$, a separate time-series forecasting model $f_{c}(\cdot;\theta_{c})$ was trained using supervised input–output pairs generated with a sliding window of size $L$. The forecaster is designed to predict a future segment of length $L$, acting as a building block to construct a complete sequence of total length $H$ (e.g., $H=500$). Training pairs were constructed as:
$(x_{t},y_{t})=\big(y^{(i)}_{t:t+L-1},\;y^{(i)}_{t+L:t+2L-1}\big)$, (1)
where $x_{t}\in\mathbb{R}^{L}$ is the context window and $y_{t}\in\mathbb{R}^{L}$ is the target segment to be predicted. For all experiments, we varied $L\in\{10,25,50,100,250\}$. Model parameters were optimized using the Huber loss [meyer2021alternative]:
@@
where $D_{c}=\{(y^{(i)},c^{(i)})\in D\mid c^{(i)}=c\}$. All models were trained with batch size 32 and a maximum of 1,000 optimization steps at a sampling rate of 100 Hz (10 ms interval).
After training, each model $f_{c}$ was used to generate synthetic EEG signals through a sliding window reconstruction strategy. Instead of relying solely on previously generated predictions, the model used temporally aligned context extracted from the original signal to maintain stable temporal structure, while generating new model-based variations at each step. Given a context window $x^{(i)}\in\mathbb{R}^{L}$ sampled from $y^{(i)}$, the model iteratively predicts non-overlapping $L$-step segments:
$\tilde{y}^{(i)}_{k}=f_{c^{(i)}}(x^{(i)}_{k};\theta_{c^{(i)}}^{\ast}),$ (3)
which are sequentially concatenated until the desired length $H$ is reached. The final synthetic dataset is defined as:
$\hat{D}=\{(\hat{y}^{(i)},c^{(i)})\}_{i=1}^{N},\qquad\hat{y}^{(i)}\in\mathbb{R}^{H}.$ (4)
Algorithm 1 summarizes the synthesis procedure.
@@
4   Construct $D_{c}=\{(x_{t},y_{t})\}$ via sliding windows
5 $\theta_{c}^{\ast}\leftarrow\arg\min_{\theta_{c}}\frac{1}{|D_{c}|}\!\sum_{(x_{t},y_{t})\in D_{c}}\mathcal{L}_{\text{Huber}}\big(f_{c}(x_{t};\theta_{c}),y_{t}\big)$
@@ 10 $\hat{y}\leftarrow$[\,]
11 for $k=1$ to $H/L$ do
12      Extract context $x^{(i)}_{k}$ of length $L$ from $y^{(i)}$
13    $\tilde{y}^{(i)}_{k}\leftarrow f_{c^{(i)}}(x^{(i)}_{k};\theta_{c^{(i)}}^{\ast})$
14      Append $\tilde{y}^{(i)}_{k}$ to $\hat{y}$
@@ 2.3 Evaluation Framework
To evaluate synthetic data utility, we designed a classification framework with three training conditions: original data only (O), synthetic data only (S), and combined original and synthetic data (O+S). Let $D_{\text{orig}}=\{(y^{(i)},c^{(i)})\}_{i=1}^{N}$ and $D_{\text{syn}}=\{(\hat{y}^{(i)},c^{(i)})\}_{i=1}^{N}$ represent the original and synthetic datasets. In the O+S condition, the training set was constructed as:
@@
Each time-series sample $u^{(i)}\in\mathbb{R}^{H}$ (representing either a real $y^{(i)}$ or synthetic $\hat{y}^{(i)}$ signal) was converted into a time-frequency representation using the short-time Fourier transform (STFT):
$S^{(i)}(f,t)=\left|\sum_{\tau=0}^{H-1}u^{(i)}(\tau)\,w(\tau-t)\,e^{-j2\pi f\tau}\right|^{2},$ (6)
@@
Model training was performed on a fixed train/test split, with training and testing sets predefined for each subject. The classifier model was a ResNet-18 [he2016deep], trained using stochastic gradient descent with a learning rate of $10^{-4}$. Note that while ResNet-18 served as the default classifier, the framework is compatible with other image classification models. To ensure reproducibility, all runs were executed with five random seeds, using AMD Instinct MI300X GPUs supported by AMD Developer Cloud credits.
@@
$\mathcal{L}(\theta)=-\sum_{i}\sum_{k=1}^{|\mathcal{C}|}\mathbbm{1}[c^{(i)}=k]\log p_{\theta}^{(i)}(k),$ (7)
where $\mathbbm{1}[\cdot]$ is the indicator function, $c^{(i)}$ is the ground-truth class label, $p_{\theta}^{(i)}(k)$ represents the predicted probability for class $k$ on sample $i$, and $|\mathcal{C}|$ denotes the number of classes. The original-only (O) and synthetic-only (S) conditions followed the same preprocessing and evaluation procedures as the O+S condition, ensuring consistency across all experimental settings.
@@ 2.4 Performance Metrics
Model performance was quantitatively assessed using accuracy and F1-score across all training conditions. Let $c^{(i)}\in\mathcal{C}$ and $\hat{c}^{(i)}\in\mathcal{C}$ denote the ground-truth and predicted class labels for the $i$-th sample in the test set. The primary evaluation metric was classification accuracy, computed as:
$\text{Accuracy}=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\mathbbm{1}[c^{(i)}=\hat{c}^{(i)}],$ (8)
where $N_{\text{test}}$ represents the total number of test samples.
To evaluate class-wise performance, we computed F1-score for each class $k\in\mathcal{C}$. Let $\text{TP}_{k}$, $\text{FP}_{k}$, and $\text{FN}_{k}$ denote the number of true positives, false positives, and false negatives for class $k$. Precision and recall were defined as:
$\text{Precision}_{k}=\frac{\text{TP}_{k}}{\text{TP}_{k}+\text{FP}_{k}},\quad\text{Recall}_{k}=\frac{\text{TP}_{k}}{\text{TP}_{k}+\text{FN}_{k}}.$ (9)
@@
$\text{F1}_{k}=2\cdot\frac{\text{Precision}_{k}\cdot\text{Recall}_{k}}{\text{Precision}_{k}+\text{Recall}_{k}}.$ (10)